OpenStudy (anonymous):

Write an equation in standard form of the line that passes through the given points. (–2, 5), (3, 5)

OpenStudy (eyust707):

oh you got this one!

OpenStudy (hoblos):

y=5

OpenStudy (anonymous):

But isn't it unidentified ?

OpenStudy (eyust707):

ok so when i was learning these i first would graph them

OpenStudy (anonymous):

I would do m = y2-y1/x2-x1

OpenStudy (eyust707):

yea!

OpenStudy (eyust707):

thats m

OpenStudy (anonymous):

then y-y1=m(x-x1)

OpenStudy (eyust707):

no remember this (x,y)

OpenStudy (eyust707):

the first number is an x component the second one is the y component

OpenStudy (anonymous):

5-5/3-(-2) = m

OpenStudy (eyust707):

correct

OpenStudy (anonymous):

0/5 or 5?

OpenStudy (eyust707):

0/5 = 0

OpenStudy (eyust707):

y=mx + b y=(0)x + b

OpenStudy (anonymous):

5 = (0)(-2)+b ? I'm using the first two points.

OpenStudy (eyust707):

so then b = ???

OpenStudy (anonymous):

5

OpenStudy (eyust707):

yep b= 5 m=0

OpenStudy (eyust707):

put those into y=mx+b

OpenStudy (anonymous):

y = x+5?

OpenStudy (eyust707):

you forgot to put in the 0

OpenStudy (anonymous):

y = 0x+5?

OpenStudy (eyust707):

yep and xero time anything equals???

OpenStudy (eyust707):

zero*

OpenStudy (anonymous):

0 !

OpenStudy (eyust707):

so the simple form is??

OpenStudy (anonymous):

m

OpenStudy (eyust707):

nooo

OpenStudy (anonymous):

:(

OpenStudy (eyust707):

y= 0x + 5 You said that 0 times x ='s 0 y = 0 + 5 y= ???

OpenStudy (anonymous):

5 !!!!!! Oops .

OpenStudy (eyust707):

y=5

OpenStudy (eyust707):

that is the correct answer! good job!

OpenStudy (anonymous):

So that's it ? Becuase I still have to write it in standard form.

OpenStudy (eyust707):